OpenStudy (anonymous):

Simplify: 8x^2 - x^3/16 - 2x

OpenStudy (anonymous):

\[\frac{ 8x ^{2}-x ^{3} }{ 16-2x}\]

OpenStudy (anonymous):

7x^2 / 2(8-x)

OpenStudy (anonymous):

Can you show me the work? @Annetta_Martin

OpenStudy (mertsj):

Can you factor x^2 out of the numerator?

OpenStudy (anonymous):

Yeah , right? @Mertsj

OpenStudy (mertsj):

Did you do it?

OpenStudy (anonymous):

I'm not good at factoring. @Mertsj

OpenStudy (anonymous):

How would you factor it out entirely? @Mertsj

thomaster (thomaster):

\[\frac{ 8x^2-x^2 }{ 16-2x }\] \[\ 8x^2-x^2 = 7x^2\] 16-2x = 2(8-x) because 2*8=16 and 2*x = 2x so: \[\frac{ 8x^2-x^2 }{ 16-2x }=\frac{ 7x^2 }{ x(8-x) }\]

OpenStudy (anonymous):

Thank you! But the equation is 8x2-x^3 not x squared. @thomaster

OpenStudy (radar):

F